# Logarithmic delocalization of low temperature 3D Ising and Potts interfaces above a hard floor

### v1

@@ Abstract.
We study the entropic repulsion of the low temperature 3D Ising and Potts interface in an $n\times n\times n$ box with blue boundary conditions on its bottom face (the hard floor), and red boundary conditions on its other five faces. For Ising, Fröhlich and Pfister proved in 1987 that the typical interface height above the origin diverges (non-quantitatively), via correlation inequalities special to the Ising model; no such result was known for Potts. We show for both the Ising and Potts models that the entropic repulsion fully overcomes the potentially attractive interaction with the floor, and obtain a logarithmically diverging lower bound on the typical interface height. This is complemented by a conjecturally sharp upper bound of $\lfloor\xi^{-1}\log n\rfloor$ where $\xi$ is the rate function for a point-to-plane non-red connection under the infinite volume red measure. The proof goes through a coupled random-cluster interface to overcome the potential attractive interaction with the boundary, and a coupled fuzzy Potts model to reduce the upper bound to a simpler setting where the repulsion is attained by conditioning a no-floor interface to lie in the upper half-space.
@@ 1. Introduction
These boundary conditions induce a $\mathsf{blue}$-to-non-$\mathsf{blue}$ interface, denoted here by $\mathcal{I}_{\mathsf{blue}}$, consisting of the maximal connected component of plaquettes separating vertices in the $\mathsf{blue}$ phase (either in the $\mathsf{blue}$ component of the boundary, or in non-$\mathsf{blue}$ bubbles encapsulated by that component), from the rest. (The plaquettes of the interface are unit squares bounded by vertices of $\mathbb{Z}^{3}$, and two plaquettes are considered adjacent if they share an edge.) Regard the height of each plaquette $f\in\mathcal{I}_{\mathsf{blue}}$ as the difference of the ${\mathfrak{e}_{3}}$-coordinate of its midpoint and the ${\mathfrak{e}_{3}}$-coordinate of the floor, which we view as having height zero, and denote it by $\operatorname{ht}(f)$. In this paper, we study the typical height of plaquettes in the bulk of the interface $\mathcal{I}_{\mathsf{blue}}$, as it balances competing forces of rigidity (which would hold at large $\beta$ if instead the boundary conditions changed color in the middle of the box), entropic repulsion away from the hard floor, and a complicated interaction between the interface and the floor through the potential finite bubbles of non-$\mathsf{blue}$ below the interface that the hard floor precludes.
@@
Rigorous quantitative results on entropic repulsion have generally been restricted to models only on interfaces (as opposed to entire spin configurations) and that are height functions (no overhangs). The model closest to the Ising/Potts interfaces is the—well-studied in its own right—Solid-On-Solid (SOS) model. There, the picture of entropic repulsion at low-temperatures is now precisely understood; the SOS interface in $\Lambda_{n}$ with 00 boundary conditions and nonnegative heights is such that all but $\varepsilon_{\beta}n^{2}$ many plaquettes of the interface are at height $\lfloor\frac{1}{4\beta}\log n\rfloor+O(1)$ [2] with more precise quantities like the shape and fluctuations of the level curve for this highest height themselves being the subject of much rich investigation [3, 1]. Entropic repulsion, and the ensuing shape of the interface above a hard floor, have also been studied extensively for related height function models like [20]; one element of all such analyses is a good understanding of the exponential rates for typical and maximal height deviations of the rigid surface before placing it near a floor.
In recent years, the precise law of height deviations, and maximum height, in the 3D Ising interface itself were derived in [11, 10], and for the random-cluster and Potts interfaces in [5]. In the Ising case, using this, a phenomenon of delocalization due to entropic repulsion was sharply studied when the repulsion is generated by what we will refer to as a soft floor, where the model has no boundary condition floor, but the interface is conditioned to stay above height zero [13]. This may feel qualitatively the same, and in the case of the height function models like SOS it is indeed the same. However, the boundary conditions at height zero is an $\exp(cn^{2})$ tilt of the measure without them, and the interactions of the interface to the boundary conditions on the floor are non-explicit and quite complicated. Moreover, since the interface is uniformly within height $O(\log n)$ of the floor, these interactions are on the same order as the entropy-induced repulsion effects and compete. We describe this and how we overcome it in more detail in Section 1.2.
@@ 1.1. Main result
Under the infinite-volume measure $\mu_{\mathbb{Z}^{3}}^{\mathsf{red}}$, at large enough $\beta$, there is almost surely an infinite connected component of $\mathsf{red}$ vertices, call it $\mathcal{V}_{\mathsf{red}}$, and all connected components of $\mathcal{V}_{\mathsf{red}}^{c}=\mathbb{Z}^{3}\setminus\mathcal{V}_{\mathsf{red}}$ are finite. Define the finite point-to-plane connectivity rate by
@@
A sub-additivity argument can be used to show that
@@
exists (in fact, we show in Lemma 4.6 that $\xi_{h}$ is approximately additive, so that $\xi_{h}=\xi h+O(1)$ for all $h$). A simple Peierls argument implies that $\xi\in[4\beta-C,4\beta+C]$ for all $h\geq 1$, for an absolute constant $C>0$. With the rate quantity $\xi$ in hand, we define
@@
Fix $q\geq 2$ and $\beta$ large enough. Consider $\mu_{\Lambda_{n}}^{\mathsf{fl}}$, the $q$-state Potts model on the $n\times n\times n$ box $\Lambda_{n}$ with boundary conditions $\mathsf{blue}$ on its bottom face and $\mathsf{red}$ elsewhere, and let $\mathcal{I}_{\mathsf{blue}}$ be the interface separating the $\mathsf{blue}$ and non-$\mathsf{blue}$ phases as defined above. There exists a sequence $\varepsilon_{\beta}\downarrow 0$ as $\beta\uparrow\infty$ such that with probability $1-o(1)$,
@@
The upper bound of Theorem 1.1 is conjecturally sharp. Indeed, for general $q\geq 2$, the height $h_{n}^{*}$ is exactly half of the maximum height oscillation $M_{n}=\max\{|\operatorname{ht}(f)|:f\in\mathcal{I}_{\mathsf{blue}}\}$ under the no-floor Dobrushin boundary condition measure $\mu_{\Lambda_{n}}^{\mathsf{dob}}$ as identified in [10, 5]. This matches the scaling relation between the repelled typical height, and no-floor maximum height in the SOS approximation [2]. To be precise, [5, Theorem 1.3] identified $M_{n}$ less explicitly in terms of a different rate, which they called the pillar rate. Part of our work in this paper (Theorem 4.1) identifies that pillar rate with $\xi_{h}$. This implies that under $\mu_{\Lambda_{n}}^{\mathsf{dob}}$, one has $M_{n}=2h_{n}^{*}+O(1)$ with probability $1-o(1)$.
@@
By the rigidity of all these interfaces, the typical height of $\mathcal{I}_{\mathsf{top}}$ (and the analogously defined random-cluster interface $\mathcal{I}_{\mathsf{bot}}$) and the typical height of the Potts interface $\mathcal{I}_{\mathsf{blue}}$ are all the same with high probability: see Lemma 2.26. As a byproduct, Theorem 1.1 also holds (with the same $h_{n}^{*}$) for the interfaces $\mathcal{I}_{\mathsf{top}},\mathcal{I}_{\mathsf{bot}}$. Perhaps surprisingly, the upper bound of $h_{n}^{*}$ proves that the typical height of $\mathcal{I}_{\mathsf{bot}}$ is not governed by the deviation rates of the no-floor $\mathcal{I}_{\mathsf{bot}}$ interface (which are known to be strictly smaller than $\xi$ by [5]) as one may initially have speculated.
@@ 1.2. Further prior work and proof ideas
Let us begin with a history of quantitative lower bounds for entropic repulsion of Ising-style interfaces that simplify some aspect to not have the delicate potential attractive interaction to the boundary of the hard-floor case. In 1991, Melicherčík [23] studied a distribution over Ising-type interfaces (allowing overhangs), but with the distribution only being over the interface, not having any finite bubbles above and below the interface, and established that the interface rises to height order $\log n$ when conditioned to be non-negative. The presence of the full spin configurations (including the possibility of finite bubbles below the interface) causes a well-known difficulty for studying Ising interfaces in the presence of nearby boundary conditions due to complicated, possibly attractive, interactions between the interface and the boundary floor. (In $d=2$ where the phenomenology is very different, the works [19, 18] required significant ideas to deal with the potential for these interactions to attract the interface to the boundary floor. The latter of these works showed first that the interface is typically at least distance $n^{\varepsilon}$ away from the floor, whence the interactions become negligible. However, in the $d=3$ case where the repulsion is only of order $\log n$, there is no hope for such an approach to controlling the interactions.) Holický and Zahradník [17] sketched in 1993 how some of the steps in [23] can be extended to incorporate the interactions in the true 3D Ising measure, postponing the details to a future version (but to our knowledge there was still no rigorous proof of this even in this Ising case). The recent work [13] proved an entropic repulsion transition for the interface in the true 3D Ising measure (including the bubbles) at height $h_{n}^{*}$ when the floor is soft, i.e., where the interface is conditioned to be above height zero. Effectively, this incorporated self-interactions of the interface through the bubbles, but still did not have to handle interactions with a nearby hard floor.
@@
Dealing with the potential attractive interactions between the interface and the hard floor is a significant part of establishing our lower bound of $(1-\varepsilon_{\beta})h_{n}^{*}$ and is achieved by instead working with a certain “top” subset $\mathcal{I}_{\mathsf{top}}$ of a coupled random-cluster model’s order-order interface (see Definition 2.1). For this specific marginal of the random-cluster interface, the interaction with the boundary is approximately in the right direction, in that it is less costly to apply maps that lift the interface in the presence of the hard floor, than in its soft-floor analogue. The reason this lower bound does not match $h_{n}^{*}$ is because the rates for downwards oscillations of $\mathcal{I}_{\mathsf{top}}$ are the source of our lower bound, but are distinct from those of $\mathcal{I}_{\mathsf{blue}}$ that we expect to govern the repulsion in $\mu_{\Lambda_{n}}^{\mathsf{fl}}$. This lower bound (and a logarithmic order upper bound) are proven in Section 2. Notably, these proofs of logarithmic order delocalization are essentially self-contained and do not use any of the more combinatorial walls, ceilings, and pillars machinery core to recent work in the subject.
@@ 2. Logarithmic upper and lower bounds
The main goal of this section is to establish the lower bound of $(1-\varepsilon_{\beta})h_{n}^{*}$ for the typical height of the plaquettes of $\mathcal{I}_{\mathsf{blue}}$. As mentioned in Section 1.2, we approach this via a coupled random-cluster model where the bottom boundary face of $\Lambda_{n}$ is conditioned not to be be connected to the five other boundary faces, inducing an interface $\mathcal{I}_{\mathsf{full}}$ of plaquettes dual to closed edges separating the two components. (By a simple energy-entropy tradeoff, when $\beta$ is large, the number of plaquettes in interfaces will not exceed $(1+\varepsilon_{\beta})n^{2}$, and so most of the plaquettes of $\mathcal{I}_{\mathsf{blue}}$ coincide with the plaquettes of $\mathcal{I}_{\mathsf{top}}$ (which bounds the open connected component of the five $\mathsf{red}$ boundary faces). Thus, lower and upper bounds on the typical height of one of these interfaces also bound the other.)
@@
    Control on repulsive interactions with hard floor via fuzzy Potts. Step (2) necessitates some non-trivial upper bound on the maximum height of the interface in the hard-floor setting. In Proposition 2.9, we use a delicate revealing and FKG argument with the fuzzy Potts representation of the model to show a monotonicity in the opposite direction of step (1) for the $\mathcal{I}_{\mathsf{blue}}$ interface: the interactions of $\mathcal{I}_{\mathsf{blue}}$ with the hard floor are only more attractive than in the soft-floor case—actually with boundary conditions flipping color at height $\mathsf{h}=C\log n$ instead of 00. In this latter setting, an $O(\log n)$ upper bound is easy to deliver. This same monotonicity is used in Section 3 to give the sharp $h_{n}^{*}$ upper bound on the height of $\mathcal{I}_{\mathsf{blue}}$.
@@ 2.1. Random-cluster, FK–Potts and fuzzy Potts models
Let $G=(V,E)$ be a finite graph. The random-cluster model on $G$, also known as the Fortuin–Kasteleyn (FK) model, is a probability distribution over subsets $\omega$ of the edge set $E$, with parameters $10<p<$ 1 and $q>0$, given by
@@
The $q$-state fuzzy Potts model [^1] The definition here is the one from [4], and the special case of the $(1,q-1)$-state fuzzy Potts model from [16]. is the pushforward measure
@@
To impose a boundary condition $\eta$ on $\Lambda_{n,m}$, one adds to the graph its outer vertex boundary in $(\mathbb{Z}+\frac{1}{2})^{3}$ to get the outer vertex boundary set $\partial\Lambda_{n,m}=\{v\in\mathbb{Z}^{d}\setminus\Lambda_{n,m}:v\sim\Lambda_{n,m}\}$, and forces those vertices to take certain colors $\eta$, e.g., $\mathsf{red}$ on boundary vertices in the upper half-space and $\mathsf{blue}$ on the lower half-space to obtain $\mu_{\Lambda_{n,m}}^{\mathsf{fl}}$. We write $\mu_{n}^{\mathsf{fl}}=\mu_{\Lambda_{n,n}}^{\mathsf{fl}}$ for brevity, denoting the analogous fuzzy Potts model by $\phi_{n}^{\mathsf{fl}}$. When a boundary condition $\eta$ is understood from context, e.g., when considering $\mu_{n}^{\mathsf{fl}}$, we use $\partial_{\mathsf{red}}\Lambda_{n,m}$ (respectively, $\partial_{\mathsf{blue}}\Lambda_{n,m}$) to denote the subsets of $\partial\Lambda_{n,m}$ getting color $\mathsf{red}$ (resp., $\mathsf{blue}$) under $\eta$.
@@
One then recovers $\sigma\sim\mu_{\Lambda_{n,m}}^{\mathsf{fl}}$ from $\omega\sim\pi^{\mathsf{w},\mathsf{w}}_{\Lambda_{n,m}}(\cdot\mid\mathfrak{D}_{n,m})$ via giving the cluster of $\partial_{\mathsf{blue}}\Lambda_{n,m}$ the color $\mathsf{blue}$, giving the cluster of $\partial_{\mathsf{red}}\Lambda_{n,m}$ the color $\mathsf{red}$, and then coloring every remaining open cluster of $\omega$ uniformly and independently (cf., e.g., [9, §2.2],[21, Fact 3.4 and Cor. 3.5] where this coupling with disconnection events was previously used). We will use the abbreviated notation $\bar{\pi}_{n}^{\mathsf{fl}}:=\pi_{\Lambda_{n,n}}^{\mathsf{w},\mathsf{w}}(\cdot\mid\mathfrak{D}_{n,n})$, and ${\mathbf{P}}_{n}^{\mathsf{fl}}$ for the joint FK–Potts distribution whose marginals are $\mu_{n}^{\mathsf{fl}}$ and $\bar{\pi}_{n}^{\mathsf{fl}}$.
@@
An important ingredient in the proofs is comparison between the hard floor setting of $\mu_{n}^{\mathsf{fl}}$ and the “soft floor” setting of the conditional distribution of $\mu_{n}^{\mathsf{h}}$ given that $\mathcal{I}_{\mathsf{blue}}$ lies above height 00, namely
@@
Let $(\sigma,\omega)\sim{\mathbf{P}}_{n}^{\mathsf{fl}}$ or $(\sigma,\omega)\sim{\mathbf{P}}_{n}^{\mathsf{h}}$.
@@
    Let $\mathcal{V}_{\mathsf{blue}}$ be the connected component (via adjacency in $(\mathbb{Z}+\frac{1}{2})^{3}$) of $\mathsf{blue}$ vertices in $\sigma$ incident to the $\mathsf{blue}$ boundary vertices; augment it to $\widehat{\mathcal{V}}_{\mathsf{blue}}$ by adding all connected components of $(\mathbb{Z}+\frac{1}{2})^{3}\setminus\mathcal{V}_{\mathsf{blue}}$ whose outer vertex boundary is a subset of $\mathcal{V}_{\mathsf{blue}}$. Define $\mathcal{V}_{\mathsf{red}}$, $\widehat{\mathcal{V}}_{\mathsf{red}}$ analogously.

### v2

@@ Abstract.
We study the entropic repulsion of the low temperature 3D Ising and Potts interface in an $n\times n\times n$ box with blue boundary conditions on its bottom face (the hard floor), and red boundary conditions on its other five faces. For Ising, Fröhlich and Pfister proved in 1987 that the typical interface height above the origin diverges (non-quantitatively), via correlation inequalities special to the Ising model; no such result was known for Potts. We show for both the Ising and Potts models that the entropic repulsion fully overcomes the potentially attractive interaction with the floor, and obtain a logarithmically diverging lower bound on the typical interface height. This is complemented by a conjecturally sharp upper bound of $\lfloor\xi^{-1}\log n\rfloor$ where $\xi$ is the rate function for a point-to-plane non-red connection under the infinite volume red measure. The proof goes through a coupled random-cluster interface to overcome the potentially attractive interaction with the boundary, and a coupled fuzzy Potts model to reduce the upper bound to a simpler setting where the repulsion is attained by conditioning a no-floor interface to lie in the upper half-space.
@@ 1. Introduction
These boundary conditions induce a $\mathsf{blue}$-to-non-$\mathsf{blue}$ interface, denoted here by $\mathcal{I}_{\mathsf{blue}}$, consisting of the set of plaquettes separating vertices in the $\mathsf{blue}$ phase (either in the $\mathsf{blue}$ component of the boundary, or in non-$\mathsf{blue}$ bubbles encapsulated by that component), from the rest (see Definition 2.1 for the formal definition). (The plaquettes of the interface are unit squares bounded by vertices of $\mathbb{Z}^{3}$, and two plaquettes are considered adjacent if they share an edge.) Regard the height of each plaquette $f\in\mathcal{I}_{\mathsf{blue}}$ as the difference of the $\mathfrak{e}_{3}$}-coordinate of its midpoint and the $\mathfrak{e}_{3}$}-coordinate of the floor, which we view as having height zero, and denote it by $\operatorname{ht}(f)$. In this paper, we study the typical height of plaquettes in the bulk of the interface $\mathcal{I}_{\mathsf{blue}}$, as it balances competing forces of rigidity (which would hold at large $\beta$ if instead the boundary conditions changed color in the middle of the box), entropic repulsion away from the hard floor, and a complicated interaction between the interface and the floor through the potential finite bubbles of non-$\mathsf{blue}$ below the interface that the hard floor precludes.
@@
Rigorous quantitative results on entropic repulsion have generally been restricted to models only on interfaces (as opposed to entire spin configurations) and that are height functions (no overhangs). The model closest to the Ising/Potts interfaces is the—well-studied in its own right—Solid-On-Solid (SOS) model. There, the picture of entropic repulsion at low-temperatures is now precisely understood; the SOS interface in $\Lambda_{n}$ with 0 boundary conditions and nonnegative heights is such that all but $\varepsilon_{\beta}n^{2}$ many plaquettes of the interface are at height $\lfloor\frac{1}{4\beta}\log n\rfloor+O(1)$ [2] with more precise quantities like the shape and fluctuations of the level curve for this highest height themselves being the subject of much rich investigation [3, 1]. Entropic repulsion, and the ensuing shape of the interface above a hard floor, have also been studied extensively for related height function models like [21]; see the survey [20] for more rich interface phenomena in SOS and related height function models. One element of all such analyses is a good understanding of the exponential rates for typical and maximal height deviations of the rigid surface before placing it near a floor.
In recent years, the precise law of height deviations, and maximum height, in the 3D Ising interface itself were derived in [11, 10], and for the random-cluster and Potts interfaces in [5]. In the Ising case, using this, a phenomenon of delocalization due to entropic repulsion was sharply studied when the repulsion is generated by what we will refer to as a soft floor, where the model has no boundary condition floor, but the interface is conditioned to stay above height zero [13]. This may feel qualitatively the same, and in the case of the height function models like SOS it is indeed the same. However, the boundary conditions at height zero are an $\exp(cn^{2})$ tilt of the measure without them, and the interactions of the interface to the boundary conditions on the floor are non-explicit and quite complicated. Moreover, since the interface is uniformly within height $O(\log n)$ of the floor, these interactions are on the same order as the entropy-induced repulsion effects and compete. In particular, these interactions could pin the surface to the floor and prevent it rising to logarithmic heights. We describe this and how we overcome it in more detail in Section 1.2.
@@ 1.1. Main result
Under the infinite-volume measure $\mu_{\mathbb{Z}^{3}}^{\mathsf{red}}$, at large enough $\beta$, there is almost surely an infinite connected component of $\mathsf{red}$ vertices, call it $\mathcal{V}_{\mathsf{red}}$, and all connected components of $\mathcal{V}_{\mathsf{red}}^{c}=\mathbb{Z}^{3}\setminus\mathcal{V}_{\mathsf{red}}$ are finite. See, e.g., [15, Chapter 7] for more on the $\beta$ large regime of the Potts model in $d\geq 3$. Define the finite point-to-plane connectivity rate by
@@
As we will see in Section 4, a sub-additivity argument using Fekete’s Lemma can show that
@@
exists (in fact, we show in Lemma 4.6 that $\xi_{h}$ is approximately additive, so that $\xi_{h}=\xi h+O(1)$ for all $h$). A simple Peierls argument (and the fact that such a non-red connection implies a finite random-cluster component of size hh)—see, e.g., [15, Thm 7.3.2]—implies that $\xi\in[4\beta-C,4\beta+C]$ for all $h\geq 1$, for an absolute constant $C>0$. With the rate quantity $\xi$ in hand, we define
@@
Fix $q\geq 2$ and $\beta$ large enough. Consider $\mu_{\Lambda_{n}}^{\mathsf{fl}}$, the $q$-state Potts model on the $n\times n\times n$ box $\Lambda_{n}$ with boundary conditions $\mathsf{blue}$ on its bottom face and $\mathsf{red}$ elsewhere, and let $\mathcal{I}_{\mathsf{blue}}$ be the interface separating the $\mathsf{blue}$ and non-$\mathsf{blue}$ phases as defined above (formally defined in Definition 2.1). There exists a sequence $\varepsilon_{\beta}\downarrow 0$ as $\beta\uparrow\infty$ such that with probability 1o(1)1-o(1),
@@
The upper bound of Theorem 1.1 is conjecturally sharp. Indeed, for general $q\geq 2$, the height $h_{n}^{*}$ is exactly half of the maximum height oscillation $M_{n}=\max\{|\operatorname{ht}(f)|:f\in\mathcal{I}_{\mathsf{blue}}\}$ under the no-floor Dobrushin boundary condition measure $\mu_{\Lambda_{n}}^{\mathsf{dob}}$ as identified in [10, 5]. This matches the scaling relation between the repelled typical height, and no-floor maximum height in the SOS approximation [2]. To be precise, [5, Theorem 1.3] identified $M_{n}$ less explicitly in terms of a different rate, which they called the pillar rate. Part of our work in this paper (Theorem 4.1) identifies that pillar rate with $\xi_{h}$. This implies that under $\mu_{\Lambda_{n}}^{\mathsf{dob}}$, one has $M_{n}=2h_{n}^{*}+O(1)$ with probability 1o(1)1-o(1).
@@
As a consequence of Lemma 2.27, with high probability the typical height of $\mathcal{I}_{\mathsf{top}}$ (and the analogously defined random-cluster interface $\mathcal{I}_{\mathsf{bot}}$), and the typical height of the Potts interface $\mathcal{I}_{\mathsf{blue}}$ are the same. As a byproduct, Theorem 1.1 also holds (with the same $h_{n}^{*}$) for the interfaces $\mathcal{I}_{\mathsf{top}},\mathcal{I}_{\mathsf{bot}}$. At the same time, the intuition behind the entropic repulsion phenomenon would suggest that the interface $\mathcal{I}_{\mathsf{bot}}$ would rise to height $\xi_{\mathsf{bot}}^{-1}\log n$, where $\xi_{\mathsf{bot}}$ is the relevant random-cluster connection rate. Perhaps surprisingly, the upper bound of $h_{n}^{*}$ proves that the typical height of $\mathcal{I}_{\mathsf{bot}}$ is $\Omega(\log n)$ smaller than this prediction (as $\xi_{\mathsf{bot}}^{-1}$ is known to be strictly smaller than $\xi$ by [5]).
@@ 1.2. Further prior work and proof ideas
Let us begin with a history of quantitative lower bounds for entropic repulsion of Ising-style interfaces that simplify some aspect to not have the delicate potentially attractive interaction to the boundary of the hard-floor case. In 1991, Melicherčík [24] studied a distribution over Ising-type interfaces (allowing overhangs), but with the distribution only being over the interface, not having any finite bubbles above and below the interface, and established that the interface rises to height order $\log n$ when conditioned to be non-negative. The presence of the full spin configurations (including the possibility of finite bubbles below the interface) causes a well-known difficulty for studying Ising interfaces in the presence of nearby boundary conditions due to complicated, possibly attractive, interactions between the interface and the boundary floor. (In $d=2$ where the phenomenology is very different, the works [19, 18] required significant ideas to deal with the potential for these interactions to attract the interface to the boundary floor. The latter of these works showed first that the interface is typically at least distance $n^{\varepsilon}$ away from the floor, whence the interactions become negligible. However, in the $d=3$ case where the repulsion is only of order $\log n$, there is no hope for such an approach to controlling the interactions.) Holický and Zahradník [17] sketched in 1993 how some of the steps in [24] can be extended to incorporate the interactions in the true 3D Ising measure, postponing the details to a future version (but to our knowledge there was still no rigorous proof of this even in this Ising case). The recent work [13] proved an entropic repulsion transition for the interface in the true 3D Ising measure (including the bubbles) at height $h_{n}^{*}$ when the floor is soft, i.e., where the interface is conditioned to be above height zero. Effectively, this incorporated self-interactions of the interface through the bubbles, but still did not have to handle interactions with a nearby hard floor.
@@
Dealing with the potentially attractive interactions between the interface and the hard floor is a significant part of establishing our lower bound of $(1-\varepsilon_{\beta})h_{n}^{*}$ and is achieved by instead working with a certain “top” subset $\mathcal{I}_{\mathsf{top}}$ of a coupled random-cluster model’s order-order interface (see Definition 2.1). For this specific marginal of the random-cluster interface, the interaction with the boundary is approximately in the right direction, in that it is less costly to apply maps that lift the interface in the presence of the hard floor, than in its soft-floor analogue. The reason this lower bound does not match $h_{n}^{*}$ is because the rates for downwards oscillations of $\mathcal{I}_{\mathsf{top}}$ are the source of our lower bound, but are distinct from those of $\mathcal{I}_{\mathsf{blue}}$ that we expect to govern the repulsion in $\mu_{\Lambda_{n}}^{\mathsf{fl}}$. This lower bound (and a logarithmic order upper bound) are proven in Section 2. Notably, these proofs of logarithmic order delocalization are essentially self-contained and do not use any of the more combinatorial walls, ceilings, and pillars machinery core to recent work in the subject.
@@ 2. Logarithmic upper and lower bounds
The main goal of this section is to establish the lower bound of $(1-\varepsilon_{\beta})h_{n}^{*}$ for the typical height of the plaquettes of $\mathcal{I}_{\mathsf{blue}}$. As mentioned in Section 1.2, we approach this via a coupled random-cluster model where the bottom boundary face of $\Lambda_{n}$ is conditioned not to be connected to the five other boundary faces, inducing an interface $\mathcal{I}_{\mathsf{full}}$ of plaquettes dual to closed edges separating the two components. (By a simple energy-entropy tradeoff, when $\beta$ is large, the number of plaquettes in interfaces will not exceed $(1+\varepsilon_{\beta})n^{2}$, and so most of the plaquettes of $\mathcal{I}_{\mathsf{blue}}$ coincide with the plaquettes of $\mathcal{I}_{\mathsf{top}}$ (which bounds the open connected component of the five $\mathsf{red}$ boundary faces). Thus, lower and upper bounds on the typical height of one of these interfaces also bound the other.)
@@
    Control on repulsive interactions with hard floor via fuzzy Potts. Step (2) necessitates some non-trivial upper bound on the maximum height of the interface in the hard-floor setting. In Proposition 2.9, we use a delicate revealing and FKG argument with the fuzzy Potts representation of the model to show a monotonicity in the opposite direction of step (1) for the $\mathcal{I}_{\mathsf{blue}}$ interface: the interactions of $\mathcal{I}_{\mathsf{blue}}$ with the hard floor are only more attractive than in the soft-floor case—actually with boundary conditions flipping color at height $\mathsf{h}=C\log n$ instead of 0. In this latter setting, an $O(\log n)$ upper bound is easy to deliver. This same monotonicity is used in Section 3 to give the sharp $h_{n}^{*}$ upper bound on the height of $\mathcal{I}_{\mathsf{blue}}$.
@@ 2.1. Random-cluster, FK–Potts and fuzzy Potts models
Let $G=(V,E)$ be a finite graph. The random-cluster model on $G$, also known as the Fortuin–Kasteleyn (FK) model, is a probability distribution over subsets $\omega$ of the edge set $E$, with parameters $0<p<1$ and $q>0$, given by
@@
The $q$-state fuzzy Potts model [^1] The definition here is the one from [4], and the special case of the (1,q1)(1,q-1)-state fuzzy Potts model from [16]. is the pushforward measure
@@
To impose a boundary condition $\eta$ on $\Lambda_{n,m}$, one adds to the graph its outer vertex boundary in $(\mathbb{Z}+\frac{1}{2})^{3}$ to get the outer vertex boundary set $\partial\Lambda_{n,m}=\{v\in(\mathbb{Z}+\frac{1}{2})^{3}\setminus\Lambda_{n,m}:v\sim\Lambda_{n,m}\}$, and forces those vertices to take certain colors $\eta$, e.g., $\mathsf{red}$ on boundary vertices in the upper half-space and $\mathsf{blue}$ on the lower half-space to obtain $\mu_{\Lambda_{n,m}}^{\mathsf{fl}}$. We write $\mu_{n}^{\mathsf{fl}}=\mu_{\Lambda_{n,n}}^{\mathsf{fl}}$ for brevity, denoting the analogous fuzzy Potts model by $\phi_{n}^{\mathsf{fl}}$. When a boundary condition $\eta$ is understood from context, e.g., when considering $\mu_{n}^{\mathsf{fl}}$, we use $\partial_{\mathsf{red}}\Lambda_{n,m}$ (respectively, $\partial_{\mathsf{blue}}\Lambda_{n,m}$) to denote the subsets of $\partial\Lambda_{n,m}$ getting color $\mathsf{red}$ (resp., $\mathsf{blue}$) under $\eta$.
@@
One then recovers $\sigma\sim\mu_{\Lambda_{n,m}}^{\mathsf{fl}}$ from $\omega\sim\pi^{\mathsf{w},\mathsf{w}}_{\Lambda_{n,m}}(\cdot\mid{\mathfrak{D}}_{n,m})$ via giving the cluster of $\partial_{\mathsf{blue}}\Lambda_{n,m}$ the color $\mathsf{blue}$, giving the cluster of $\partial_{\mathsf{red}}\Lambda_{n,m}$ the color $\mathsf{red}$, and then coloring every remaining open cluster of $\omega$ uniformly and independently (cf., e.g., [9, §2.2],[22, Fact 3.4 and Cor. 3.5] where this coupling with disconnection events was previously used). We will use the abbreviated notation $\bar{\pi}_{n}^{\mathsf{fl}}:=\pi_{\Lambda_{n,n}}^{\mathsf{w},\mathsf{w}}(\cdot\mid{\mathfrak{D}}_{n,n})$, and ${\mathbf{P}}_{n}^{\mathsf{fl}}$ for the joint FK–Potts distribution whose marginals are $\mu_{n}^{\mathsf{fl}}$ and $\bar{\pi}_{n}^{\mathsf{fl}}$.
@@
Refer to captionRefer to captionRefer to caption
Figure 1. A Potts configuration (left) and its two Potts interfaces $\mathcal{I}_{\mathsf{blue}}$ and $\mathcal{I}_{\mathsf{red}}$ (right).
Refer to captionRefer to captionRefer to caption
Figure 2. A random-cluster configuration (left), and its interfaces $\mathcal{I}_{\mathsf{bot}}$ and $\mathcal{I}_{\mathsf{top}}$ (right).
Figure 3. The four coupled interfaces, $\mathcal{I}_{\mathsf{top}}$ (pink highlighted), $\mathcal{I}_{\mathsf{red}}$ (red line), $\mathcal{I}_{\mathsf{blue}}$ (blue line) and $\mathcal{I}_{\mathsf{bot}}$ (cyan highlighted), showing the ordering $\mathcal{I}_{\mathsf{top}}\succeq\mathcal{I}_{\mathsf{red}}\succeq\mathcal{I}_{\mathsf{blue}}\succeq\mathcal{I}_{\mathsf{bot}}$.
@@
Let $(\sigma,\omega)\sim{\mathbf{P}}_{n}^{\mathsf{fl}}$ or $(\sigma,\omega)\sim{\mathbf{P}}_{n}^{\mathsf{h}}$. The below definition applies to both, with the sets $\mathsf{blue}$ boundary vertices”, and $\mathsf{red}$ boundary vertices” interpreted according to the setting.
@@
    Let $\mathcal{V}_{\mathsf{blue}}$ be the connected component (via adjacency in $(\mathbb{Z}+\frac{1}{2})^{3}$) of $\mathsf{blue}$ vertices in $\sigma$ incident to the $\mathsf{blue}$ boundary vertices; augment it to $\widehat{\mathcal{V}}_{\mathsf{blue}}$ by adding all connected components of $(\mathbb{Z}+\frac{1}{2})^{3}\setminus\mathcal{V}_{\mathsf{blue}}$ whose outer vertex boundary is a subset of $\mathcal{V}_{\mathsf{blue}}$. (This guarantees that all bubbles in the $\mathsf{blue}$ phase are in $\widehat{\mathcal{V}}_{\mathsf{blue}}$.) Define $\mathcal{V}_{\mathsf{red}}$, $\widehat{\mathcal{V}}_{\mathsf{red}}$ analogously. (See Fig. 1.)